Reports 1-1 of 1 Clear search Modify search
AdV-TCS (TCS studies actuation (simulations + experiments))
capocasa, degallaix, rocchi, spinicelli, taranto - 0:37 Thursday 19 May 2022 (55900) Print this report
TCS - CO2 Actuator tuning keeping the arm matching

The goal of the shift was to tune the DAS at the NI in order to compensate the CP converging lens on the North arm.

We started from the configuration prepared during the morning shift, that is the matching of the West Arm with the TCS off.

Then, Claudia switched on DAS on NI with the following parameter:
15:25 UTC: NI_DAS on with 0.1 W inner and 2 W outer

The thermalization is supposed to be longer than 2h, we then left the scan of TEM2 going on for this time. At the end of this period, we found a huge misalignment that probably spoiled the last  measurements. 
The analysis of the first DAS point (fig. 1) shows that we have an improvement in the first part followed by a degradation of the matching. Anyway, the north arm couldn't reach the same level of matching as the west one. 

We then decided to take a scan of the FSR/TEM2 before and after a realignment of the arms in order to verify if there was a difference in the estimated mismatch. 
In fig.2 we see that the realignment  improved the matching in the last points.

From now on, we block the scan and realign each ~20 minutes to avoid this problem. In agreement with Alessio, we decided to look at another DAS point by changing only one actuator. We thus increased only the outer ring power:

1336931077  18 May 2022 17:44:19 UTC  (18 May 2022 19:44:19 CEST):  doubling outer das (from ~2W to ~4W)

During the thermal transient, we unlocked and realigne few times:
1336932459  18 May 2022 18:07:21 UTC  (18 May 2022 20:07:21 CEST)
1336934045  18 May 2022 18:33:47 UTC    (18 May 2022 20:33:47 CEST)
1336935289  18 May 2022 18:54:31 UTC  (18 May 2022 20:54:31 CEST)

As the matching seems to go into the wrong direction (fig. 3), we decided to reduce the power instead.

19:02 UTC: outer ring from 4 W to 0.95 W

During the thermal transient, we unlocked and realigne few times:
1336936797  18 May 2022 19:19:39 UTC  (18 May 2022 21:19:39 CEST)
1336937565  18 May 2022 19:32:27 UTC  (18 May 2022 21:32:27 CEST)      (unwanted => green unlock?)
1336939657  18 May 2022 20:07:19 UTC  (18 May 2022 22:07:19 CEST)

The matching indeed improved (fig.4).

We continued switching off completely the outer ring:
1336940042  18 May 2022 20:13:44 UTC  (18 May 2022 22:13:44 CEST)

During the thermal transient, we unlocked and realigne few times:
1336941248  18 May 2022 20:33:50 UTC  (18 May 2022 22:33:50 CEST)
1336942679  18 May 2022 20:57:41 UTC  (18 May 2022 22:57:41 CEST)

Apparently, this step didn't improve the matching (fig5, fig 6).

A summary of the different steps is shown in figure 7.

The DAS tuning will continue tomorrow.

We leave the arm locked on IR and the only WI inner DAS ON @0.1W.

Images attached to this report
Search Help